HW help.
4. Solve the initial value problem. 2 -1 6 x(0)
Answer #1

& 4 1 6 2 ie 20 & 402 dis zlo) given initial value problem is, x 40)= (6) o 1x, los )-40) dae ( ) The characteristic equationfrom which we at once find that d, ond d must satists, 2d, +42 4dg -d, +6% = 4 d₂ lie -22, + 4 dy = 0 - di +222 = 0 On d = 22[() - (19) - () . - 2 1)(2) = (3) From the we at Once find sotisty that P. and be 쟈, 22 = 1.. A triral simple solut of thisNow initial condition 31 (0) (0) 2,(0) ( O) = 6 ) 71, 10) = -1, 7200= 6 Then. by we have. 2,00) 2c, & q = -1 210) 4, + < = 6